### v1

@@ 3.4.1 Performance benchmarking with current state of the art clustering techniques
For high-dimensional clustering analysis, there are certain popular choice of working algorithms with varying implementation techniques. Methods like DBSCAN (ester1996density) and HDBSCAN (campello2013density) are based on density of the data distributed over dimensional space, where data points in the low-density region are identified as noise/outliers and form clusters using core points. Shared nearest neighbours based method (ertoz2002new) works by grouping data points with higher similarity of shared neighbours across a k-nearest neighbour graph. bouveyron2007high proposed a high dimensional model based clustering that implements dimensionality reduction for each cluster with parsimonious covariance structure and estimate the Gaussian mixture model parameters using EM iterations. Another class of popular methods include graph based algorithms, like Leiden (an upgrade of Louvain) (traag2019louvain) that constructs a k-nearest neighbour graph from the data and detects communities (clusters) by maximizing graph modularity. Classical $K$-means (mcqueen1967some), on the other hand, is a partitioning based method that minimizes the within-cluster variance by iteratively assigning data points to their nearest centroids.
For performance on the Leukemia data set, we implemented the methods in R and chose the key parameters involved with these methods based on internal validation metrics specific to each method (details in Supplementary material). Based on corresponding optimal parameters, these methods are compared with Sparse DPMM implementation with empirical Bayes hyper-parameters, and we validated their performances with run-time (in seconds), estimated cluster numbers and ajusted Rand index (ARI). Due to diversity in methodology of these techniques, we avoided any internal metric for comparison (due to possible bias for particular technique(s)) and focused on external validation metrics. Figure 7 shows that Sparse DPMM is significantly slower ($\sim 8.091$ seconds per iteration) in comparison to alternative methods. However, the performance metrics based on estimated number of clusters and ARI scores validate the (relative) superiority of Sparse DPMM.
@@ 4 Discussion
We developed a Bayesian non-parametric mixture model using collapsed variational inference (VI) to perform unsupervised clustering. Our approach hierarchically incorporates the Dirichlet process (DP) concentration parameter $\alpha$ as a variable in the DP mixture model (DPMM) with Gaussian kernels. Based on the importance of adaptive estimation of $\alpha$ (ascolani2023clustering), the proposed framework enhances the consistency of collapsed VI model developed by kurihara2007collapsed. We also illustrated the effect of prior for covariance matrix, and concluded that a cluster-specific Sparse DPMM is the best choice. Implementing our theoretical work into an R-package vimixr, the user is allowed to choose the model and prior choice based on parameterisations given in Table 1.
@@
The Sparse DPMM provides significantly faster convergence than an MCMC slice sampling approach (hejblum2019sequential). Although the convergence speed varies with the dimension $d$ due to cluster-specific covariance structure, the Sparse DPMM performs robustly well for a large range of samples $N$, dimensions $d$ and true clusters $K_{true}$ present in the data. Comparing with conventional clustering techniques, the iterative nature of Sparse DPMM, added to the model structure complexity, leads to a slower convergence than the conventional techniques considered. However, it provides better clustering estimates with empirical Bayes hyper-parameters using gene expression data for leukemia sub-types (armstrong2002mll). Even with weaker hyper-parameter values, Sparse DPMM estimates are consistent with gene expression profiles, hence providing meaningful sub-clusters of the already known sub-types.
@@
A natural extension of Sparse DPMM is to generalize beyond Gaussian kernels for the DP base distribution parameter $G_{0}$, which becomes eminent for real-world data. For instance, overdispersed count data is better modelled using Negative Binomial distributions (fernandez2022comparison). So, incorporating Negative Binomial kernels can aid in avoiding model misspecification for clustering such data sets like gene expression data (anders2010differential). Another aspect is to introduce probabilistic feature selection variables that could be updated simultaneously (tadesse2005bayesian) or guide the clustering process itself (rouanet2024bayesian). This can provide insights on the features defining the estimated clusters and prove beneficial, both computationally and inferentially, for high-dimensional data sets.
@@
As an optimisation technique, VI algorithms heavily depend on the initial choice of latent variables (latent allocation vectors $\{z_{n}\}$ in our case) for a given data. We advocate random initialisation for these parameters, and the choice is based on variational log likelihood (VLL) values. For VI implementation, alternative approaches (zhang2018advances) can be explored for further refinement. We have also shown that the hyper-parameters’ choice for $G_{0}$ significantly effect the cluster estimation. Based on the analytical properties of VLL, empirical Bayes estimates for the hyper-parameters prove to be efficient for robust estimation. Due to their implied significance over effective dimensions and sample size, it becomes important to consider hierarchical hyper-priors for the hyper-parameter $a_{0}=b_{0}$. However, the conjugate prior for both unknown shape and rate parameters of a Gamma distribution does not have a closed form (miller1980bayesian). So, incorporating an $a_{0}=b_{0}$ hyper-prior would add another layer of analytical complexity, along with increased computational cost due to additional model parameters.
@@
The exchangeability of mixture components in DPMM introduces label switching between atoms as they are ordered sequentially through the stick-breaking construction (papaspiliopoulos2008retrospective; hastie2015sampling). Following kurihara2007collapsed, we addressed the issue by ordering the latent allocation probability updates such that the clusters are arranged in decreasing proportions. Alternatively, fruhwirth2011dealing shows identification methods for finite mixture model, that could be extrapolated to non-parametric mixtures.
5 Conflicts of interest
The authors declare that they have no competing interests.
6 Funding
This work is supported by funds from the Programmes et équipements prioritaires de recherche (PEPR) Santé Numérique SMATCH Task 2.N1, under the Agence nationale de la recherche (ANR), France.
7 Software and Data availability
Software in the form of R package vimixr with complete documentation is available on CRAN (https://cran.r-project.org/web/packages/vimixr/index.html). The Leukemia sub-type gene expression data is publicly available online at https://schlieplab.org/Static/Supplements/CompCancer/Affymetrix/armstrong-2002-v2/. The package implementation is provided in the Zenodo repository https://doi.org/10.5281/zenodo.18405688, along with metadata and results.
8 Supplementary Material
Supplementary material is available online at http://biostatistics.oxfordjournals.org.
@@ Acknowledgments
The authors thank the anonymous reviewers for their valuable suggestions. This work was supported by a French government grant managed by the Agence Nationale de la Recherche (ANR) under the France 2030 program (reference SMATCH ANR-22-PESN-0003). For heavy computation, the study used Cluster Curta, a project by Mésocentre de Calcul Intensif Aquitain (MCIA), France (https://redmine.mcia.fr/projects/cluster-curta)

### v2

@@ 3.3.1 Performance benchmarking with current state of the art clustering techniques
For high-dimensional clustering analysis, there are certain popular choice of working algorithms with varying implementation techniques. Methods like DBSCAN (Ester et al., 1996) and HDBSCAN (Campello et al., 2013) are based on density of the data distributed over dimensional space, where data points in the low-density region are identified as noise/outliers and form clusters using core points. Shared nearest neighbours based method (Ertoz et al., 2002) works by grouping data points with higher similarity of shared neighbours across a k-nearest neighbour graph. Bouveyron et al. (2007) proposed a high dimensional model based clustering that implements dimensionality reduction for each cluster with parsimonious covariance structure and estimate the Gaussian mixture model parameters using EM iterations. Another class of popular methods include graph based algorithms, like Leiden (an upgrade of Louvain, Traag et al. (2019)) that constructs a k-nearest neighbour graph from the data and detects communities (clusters) by maximizing graph modularity. Classical $K$-means (MacQueen, 1967), on the other hand, is a partitioning based method that minimizes the within-cluster variance by iteratively assigning data points to their nearest centroids.
@@
For performance on the Leukemia dataset, we implemented the methods in R and chose the key parameters involved with these methods based on internal validation metrics specific to each method (details in Supplementary material). Based on corresponding optimal parameters, these methods are compared with Sparse DPMM implementation with empirical Bayes hyper-parameters, and we validated their performances with run-time (in seconds), estimated cluster numbers and ajusted Rand index (ARI). Due to diversity in methodology of these techniques, we avoided any internal metric for comparison across methods and focused on external validation metrics. Figure 5 shows that Sparse DPMM is significantly slower ($\sim 8.091$ seconds per iteration) in comparison to alternative methods. However, the performance metrics based on estimated number of clusters and ARI scores validate the (relative) superiority of Sparse DPMM.
@@ 4 Discussion
We developed a Bayesian non-parametric mixture model using collapsed variational inference (VI) to perform unsupervised clustering. Our approach hierarchically incorporates the Dirichlet process (DP) concentration parameter $\alpha$ as a variable in the DP mixture model (DPMM) with Gaussian kernels. Based on the importance of adaptive estimation of $\alpha$ (Ascolani et al., 2023), the proposed framework enhances the consistency of collapsed VI model developed by Kurihara et al. (2007). We also illustrated the effect of prior for covariance matrix, and concluded that a cluster-specific Sparse DPMM is the best choice. Implementing our theoretical work into an R-package vimixr, we provide a probabilistic tool to perform exploratory analysis of unstructured and high-dimensional biological data with significant potential applicability.
@@
The Sparse DPMM provides significantly faster convergence than an MCMC slice sampling approach (Hejblum et al., 2019). Comparing with conventional clustering techniques, the iterative nature of Sparse DPMM, added to the model structure complexity, leads to a slower convergence. However, it provides better clustering estimates with empirical Bayes hyper-parameters using gene expression data for leukemia sub-types (Armstrong et al., 2002). Even with weaker hyper-parameters, Sparse DPMM estimates are consistent with gene expression profiles, hence providing a biologically meaningful sub-cluster of the known sub-types.
@@
A natural extension of Sparse DPMM is to generalize beyond Gaussian kernels for $G_{0}$, which becomes eminent for real-world data. For instance, overdispersed count data is better modelled using Negative Binomial distributions, thus avoiding model misspecification for clustering such datasets like gene expression data (Anders and Huber, 2010). Another aspect is to introduce probabilistic feature selection variables that could be updated simultaneously (Tadesse et al., 2005) or guide the clustering process itself (Rouanet et al., 2024). This can provide insights on the features defining the estimated clusters and prove beneficial for high-dimensional datasets.
@@
As an optimization technique, VI algorithms heavily depend on the initial choice of latent variables (latent allocation vectors $\{z_{n}\}$ in our case) for a given data. We advocate random initialization for these parameters, and the choice is based on variational log likelihood (VLL) values. For VI implementation, alternative approaches (Blei et al., 2017) can be explored for further refinement. Based on the analytical properties of VLL, empirical Bayes estimates for the $G_{0}$ hyper-parameters prove to be efficient for robust estimation. Due to their implied significance over effective dimensions and sample size, it becomes important to consider hierarchical hyper-priors for the hyper-parameter $a_{0}=b_{0}$. However, the conjugate prior for both unknown shape and rate parameters of a Gamma distribution does not have a closed form (Miller, 1980). So, incorporating an $a_{0}=b_{0}$ hyper-prior would add another layer of analytical complexity, along with increased computational cost due to additional model parameters.
@@
The exchangeability of mixture components in DPMM introduces label switching between atoms as they are ordered sequentially through the stick-breaking construction (Papaspiliopoulos and Roberts, 2008). Following Kurihara et al. (2007), we addressed the issue by ordering the latent allocation probability updates such that the clusters are arranged in decreasing proportions. Alternatively, Frühwirth-Schnatter (2011) shows identification methods for finite mixture model, that can be extrapolated to non-parametric mixtures.
@@ Acknowledgments
This work was supported by the University of Bordeaux’s Digital Public Health Graduate School, financed by the PIA 3 through the Agence Nationale de la Recherche (Ref: 17-EURE-0019), by the Programme et Equipement Prioritaire de Recheche Santé Numérique (PEPR SN) project SMATCH (Ref: 22-PESN-0003). Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour.
Supplementary Material
Supplementary material is available with this paper at the Biometrics website on Wiley Online Library.
Appendix A
A.1 Software and Data availability
Software in the form of R package vimixr with complete documentation is available on CRAN (https://cran.r-project.org/web/packages/vimixr/index.html). The Leukemia sub-type gene expression data is publicly available online at
https://schlieplab.org/Static/Supplements/CompCancer/Affymetrix/armstrong-2002-v2/. The package implementation is provided in the Zenodo repository
https://doi.org/10.5281/zenodo.18405687, along with metadata and results.